# State-Covering Trajectory Stitching for Diffusion Planners

### v2

@@ Appendix D Baseline Performance Sources
Performance scores reported for offline goal-conditioned reinforcement learning (GCRL) methods, including Goal-Conditioned Implicit Q-Learning (GCIQL) Kostrikov et al. (2022), Quasimetric RL (QRL) Wang et al. (2023), Contrastive RL (CRL) Eysenbach et al. (2022), and Hierarchical Implicit Q-Learning (HIQL) Park et al. (2023a), are sourced from Table 2 in Park et al. (2024a). Scores for diffusion-based generative planning methods explicitly designed for long-horizon generalization, including Generative Skill Chaining (GSC) Mishra et al. (2023) and Compositional Diffuser (CD) Luo et al. (2025), are sourced from Tables 1 and 2 in Luo et al. (2025).

### v3

@@ Appendix D Baseline Performance Sources
Performance scores reported for offline goal-conditioned reinforcement learning (GCRL) methods, including Goal-Conditioned Implicit Q-Learning (GCIQL) Kostrikov et al. (2022), Quasimetric RL (QRL) Wang et al. (2023), Contrastive RL (CRL) Eysenbach et al. (2022), and Hierarchical Implicit Q-Learning (HIQL) Park et al. (2023), are sourced from Table 2 in Park et al. (2025). Scores for diffusion-based generative planning methods explicitly designed for long-horizon generalization, including Generative Skill Chaining (GSC) Mishra et al. (2023) and Compositional Diffuser (CD) Luo et al. (2025), are sourced from Tables 1 and 2 in Luo et al. (2025).